OpenStudy (anonymous):

If 2x^2 - 2x - 1 = 0 then x = .. ? What....

OpenStudy (anonymous):

factor!

OpenStudy (anonymous):

duh; i dont know how to with the 2x^2 there..

OpenStudy (anonymous):

x = 1/2 (1-sqrt(3)) x = 1/2 (1+sqrt(3)) Would you like an explanation?

OpenStudy (anonymous):

Yes please!